Please help, it's a Maths question for homework needed to be handed in tomorrow.
:
A man is 8 times as old as his son. In 5 years he will be only 4 times as old. How old is the son now?
Thanks in advance :)

Answer by scott8148(6628)

"A man is 8 times as old as his son" ___ m = 8s

"In 5 years he will be only 4 times as old" ___ m + 5 = 4(s + 5)

substituting the equivalent value of m from the 1st equation into the 2nd equation ___ (8s) + 5 = 4s + 20

4s = 15


the answer is not a whole number, which probably means that the numbers in the question need to be changed